OpenStudy (anonymous):

A 2.0-kg solid cylinder of radius 0.5 m rotates at a rate of 40 rad/s about its cylindrical axis. What power is required to bring the cylinder to rest in 10 s?

OpenStudy (anonymous):

\[P = \frac{W}{t}\] \[ \quad W = \Delta KE\] \[ \qquad KE_{rot} = \frac{1}{2}I\omega^2\] \[\ \ \ \qquad I_{cyl} = \frac{1}{2}MR^2 \] \[P = \frac{\frac{1}{2} \left( \frac{1}{2}MR^2 \right) \omega^2}{t} \]

OpenStudy (anonymous):

Nice.